OpenStudy (anonymous):

A point H on a segment with endpoints B(3, -1) and Z(12, 5) partitions the segment in a 5:1 ratio. Find H. I suck at the ratios of a line :/

OpenStudy (welshfella):

there is a formula you can use (x ,y) = (x1 + k(x2-x1)) , ( y1 + k(y2 - y1)) where k = 5 / (5 + 1)

OpenStudy (mathmale):

Sketching this situation would very likely help you to estimate what the correct answer is. Would you try out the Draw utility for this purpose? Plot the 2 giv en points and then draw a line connecting them. How long is this line? Use the Pythagorean Theorem or distance formula to answer that question.

OpenStudy (anonymous):

So k is the ratio?

OpenStudy (anonymous):

And the distance of the line is 10.82

OpenStudy (anonymous):

I dont think i did it roght because i got 34.5 for x and -1 for y im gonna try again but i dont know

OpenStudy (anonymous):

I redid it and got 9 and 2/3 for x

OpenStudy (anonymous):

And i cant just estimate the right answer i have to write it in sentences with an explanation :/

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

x1 = 3 , y1 = -1 x2 = 12 , y2 = 5

OpenStudy (welshfella):

k = 5/6 so we have (3 + (5/6)( 12-3) , (-1 + (5/6)(5 -(-1))

OpenStudy (anonymous):

Okay i see where i messed up now so H should be (10.5, 4) right?

OpenStudy (welshfella):

yes